The period of a particle in SHM is 8s. at t=0, it is at te mean position. the ratio of the distances travelled by it in 1st and 2nd second is.....................
 

Grade:12

2 Answers

Arun
6 years ago
Dear Student
 
here is my detailed solution
 
2 -1) \sqrtlet the equation of motion is X =Asin(2px/8) in first second X1 =Asin(p/4) in 2 sec X2 =Asin(p/2) so 2nd sec distance covered = X2 -X1 x1 /(X2 -X1) = Asin(p/4)/[Asin(p/2) -Asin(p/4)] =1/(
 
regards
arun
Sougat Das
6 years ago
time period is 8s......so angular velocity=2π/8=π/4let y` be the distance in 1st sec.y`=a sin(π/4)let y" be the distance travelled in 2nd sec.y"=a(sin π/2-sin π/4)so y`/y"=1/(√2-1)